# GreekBarBench: A Challenging Benchmark for Free-Text Legal Reasoning and Citations

### v2

@@ Abstract
We introduce GreekBarBench, a benchmark that evaluates LLMs on legal questions across five different legal areas from the Greek Bar exams, requiring citations to statutory articles and case facts. To tackle the challenges of free-text evaluation, we propose a three-dimensional scoring system combined with an LLM-as-a-judge approach. We also develop a meta-evaluation benchmark to assess the correlation between LLM-judges and human expert evaluations, revealing that simple, span-based rubrics improve their alignment. Our systematic evaluation of $13$ proprietary and open-weight LLMs shows that even though the best models outperform average expert scores, they fall short of the 95th percentile of experts.
@@
Odysseas S. Chlapanis1,2Dimitrios Galanis3Nikolaos Aletras4Ion Androutsopoulos1,2 1Department of Informatics, Athens University of Economics and Business, Greece 2Archimedes, Athena Research Center, Greece 3Athena Research Center, Greece 4University of Sheffield, United Kingdom
@@ 1 Introduction
As legal AI assistants become increasingly prevalent, the need for realistic legal LLM benchmarks has never been more imperative. [^1] https://www.abajournal.com/web/article/aba-tech-report-finds-that-ai-adoption-is-growing-but-some-are-hesitant Most widely used legal Natural Language Processing (NLP) benchmarks (Chalkidis et al., 2022; Niklaus et al., 2023) focus on classification tasks, e.g., legal judgement prediction (Aletras et al., 2016), which have been criticized (Medvedeva and Mcbride, 2023) for being more constrained and less representative than real-world tasks. Even more recent LLM-focused legal benchmarks (Guha et al., 2023; Fei et al., 2024; Joshi et al., 2024) do not go beyond closed-form questions (e.g., multiple-choice questions), failing to capture the true complexity of legal reasoning in practice, which involves identifying, analyzing and synthesizing relevant information to reach a conclusion. Unfortunately, most existing benchmarks with challenging legal questions and free-text responses are proprietary and thus inaccessible to the research community. [^2] https://www.vals.ai/benchmarks
@@
$[2]$ Ioannis’s assistant, Penelope, accidentally applied pure acetic acid due to a mislabeled bottle, causing burns on Antonis’s face.
$[3]$ He needed plastic surgery costing €2,500 and is now seeking these costs plus €75,000 for moral damages.
@@
actions $[2]$(Civil Code 922)and Penelope is directly
liable for her negligence $[2]$(Civil Code 914).Hence,
both are liable and must compensate Antonis.
Table 1: Cropped example (English translation) from GreekBarBench. The answer requires multi-hop reasoning and citing legal articles and case facts. The spans corresponding to the scoring dimensions are highlighted in color: Facts (green), Cited Articles (blue) and Analysis (orange). Important spans are marked in bold and cited facts are denoted by square brackets. The complete example is presented in Appendix C.
@@
  • A systematic evaluation of 13 frontier and open-weight LLMs on GreekBarBench, using the best LLM-judge at GBB-JME. Top models surpass average expert performance, but not the 95th percentile of experts.
@@ 2 GreekBarBench (GBB)
2.1 Greek Bar Exams
Law graduates in Greece must pass the Greek Bar exam to become licensed attorneys. The exam evaluates candidates through practical legal questions across five key areas of law: Civil Law, Criminal Law, Commercial Law, Public Law, Lawyers’ Code. Greece’s legal system is statutory, meaning that laws are derived from legal code documents (statutes), rather than from judicial precedents (case law). The exams are open-book; candidates have access to legal code documents and are expected to cite statutory articles from them in their answers.
The available documents include: the Civil Code and Civil Procedure Code, the Criminal Code and Criminal Procedure Code, eight Commercial Law codes, eleven Public Law codes, as well as the Lawyers’ Code and the Code of Ethics for Legal Practice (see Table 3). Candidate lawyers typically approach the exam by first studying the case facts to identify the relevant legal issues. They then navigate the legal code documents to find the relevant chapter and pinpoint the exact statutory article within it to cite in support of their arguments.
@@ 2.2 Task
Benchmark Lang Citations Multi-Hop Free-Text Judge-Eval
@@
Table 2: Comparison of legal benchmarks. GreekBarBench uniquely encompasses all challenging features essential for evaluation in realistic and practical scenarios. ‘Lang’: language of dataset. ‘Citations’: legal articles. ‘Multi-hop’: reasoning using multiple sources. ‘Free-text’: open-ended responses. ‘Judge-Eval’: separate benchmark to compare LLM-judges (‘-’ means that the benchmark exists but is not publicly available).
Each instance in GreekBarBench is taken from Greek Bar exam papers. The input consists of (1) the case facts, (2) the legal question, and (3) a collection of potentially relevant chapters of statutory articles. The desired output is the correct free-text answer to a legal question, providing an analysis with citations to the case facts and the applicable statutory articles. The primary challenges are to discern significant facts, to identify the applicable articles and, finally, to analyze the outcomes of the application of the articles to answer the question (see the example in Table 1). These attributes make GreekBarBench unique compared to other legal benchmarks (Table 2).
@@ 2.3 Dataset Statistics
We collect a total of 65 exam papers; 13 exam papers from each of the five aforementioned areas. The papers and suggested solutions are publicly available in a booklet in PDF format, spanning from 2015 to 2024.444The booklet is available at https://www.lawspot.gr/nomika-nea/panellinios-diagonismos-ypopsifion-dikigoron-themata-exetaseon-kai-endeiktikes-3. The authors permit its distribution for academic research only. The booklet is converted to text format and further processed (§ 2.5) to prepare the dataset. Each exam paper includes 4.7 questions on average, resulting in a total of 310 samples. We keep the questions from 2024 (22 in total) as a semi-private test set, to avoid data contamination. [^5] With ‘semi-private’ we mean that the test set is not public, but the raw data sources are available. The remaining 288 samples comprise the public test set. The semi-private set will be updated each year with two more exam papers from each legal area (there are two examinations per year), and made publicly available.
Answering the exam questions requires citing articles from 25 legal code documents, obtained from the official website of the Greek government. [^6] https://search.et.gr/el/ Detailed statistics for these documents are presented in Table 3 per legal area. Articles are cited 931 times in total, across all exam questions. The articles within each legal code document are grouped thematically into chapters. The total number of citable articles is more than 16 thousand.
Law Areas Samples Legal Total Cited Context
@@ 2.4 Relevant Legal Context
As mentioned in Section 2.1, the Greek Bar exams are open-book, allowing candidate lawyers to navigate legal code documents to identify relevant statutory articles for the presented case. Simulating this setup presents several challenges. One approach would be implementing a standard Retrieval-Augmented Generation (RAG) pipeline, using sparse (e.g., BM25) or dense retrievers (Karpukhin et al., 2020) to select the $k$ most ‘relevant’ articles for inclusion in the LLM’s input. However, this approach suffers from three significant limitations: a) candidate lawyers taking the exams do not have access to such retrieval tools, making direct comparisons with human performance problematic; b) retrievers are prone to errors, creating a substantial risk that even with large values of $k$, the ground truth articles might not appear among the top retrieved articles; and c) as demonstrated by Krishna et al. (2025), benchmarking RAG systems requires testing multiple configurations with varying values of $k$ and, ideally, different retriever models, complicating fast integration of new LLMs.
Instead, we adopt a simplified yet challenging approach. For each legal case, we automatically collect all the ground truth articles cited across all questions pertaining to that case (using regular expressions), and identify the chapters containing the articles. We hypothesize that all the articles of the identified chapters have substantial relevance for all questions of the particular case. Therefore, we provide the entire text (all articles) of the identified chapters as the relevant legal context in the LLM input for every question of the particular case. Hence, for each question, the input context contains all the ground truth articles of the question, along with all the other articles from their chapters, as particularly close distractors. It also includes additional distractor articles from other chapters related to other questions of the same case, which increase the difficulty of citing the correct articles.
The total length of the legal context fed to the LLM per question is 62k tokens on average (Table 3). This makes the task manageable for recent LLMs with context windows exceeding 100k tokens, while still presenting a significant challenge, as LLMs often struggle to extract crucial information from extended contexts (Liu et al., 2024).
@@ 2.5 Fact Segmentation
To enhance the evaluation process (both manual and automatic), we require citations to facts in candidate answers, though this is not mandatory in official exams. Fact-citations help legal professionals in practical applications by enabling more efficient verification of answer accuracy. To facilitate fact-citation, we present the case facts as a numbered list of sentences (as shown in Table 1). For segmentation, we employ the Segment-Any-Text neural model (Frohmann et al., 2024).
@@ 2.6 Three-Dimensional Scoring System
The official evaluation committee of the Greek Bar Exams grades candidate answers on a scale of 1 to 10. Without explicit guidelines, they grade by comparing answers to the ground truth based on their discretion. Drawing inspiration from established legal research and evaluation practices (Clark and DeSanctis, 2013), and guided by our legal expert annotators (§ 5.2), we develop a novel three-dimensional scoring system to improve the evaluation process for the benchmark. The proposed approach assesses legal reasoning across three dimensions: the Facts, the Cited Articles, and the Analysis. Each dimension is rated on a scale of 1 to 10, and the final score is the average of these three. This system allows explainability through detection of specific shortcomings in the reasoning abilities of LLMs. The Facts score measures understanding of case facts; the Cited Articles evaluates the accuracy and interpretation of cited legal articles; and the Analysis evaluates the ability to apply legal articles to the facts and reach conclusions. For instance, a low Facts score indicates hallucinations, a low Cited Articles score shows difficulty in identifying applicable articles, and a low Analysis score reveals weakness in legal reasoning.
@@ 3.1 Simple LLM-Judge
As an initial approach, we designed a straightforward prompt for a simple LLM-judge. The prompt outlines the evaluation task and explicitly defines the criteria for the Facts, Cited Articles, and Analysis scores. All necessary contextual information is provided; the facts of the case, the specific legal question, the ground truth answer with the cited articles and the candidate answer to be evaluated. This context mirrors the information provided to the human annotators for the manual evaluations (§ 5.2). The required output format is clearly specified: the model must provide an explanation for each score, followed by the integer score. The complete prompt is presented in Appendix B (Fig. 5).
@@ 3.2 Span LLM-Judge
According to Clark and DeSanctis (2013), rubrics, i.e., instructions that break down an assignment in identifiable components, can significantly improve the consistency of legal writing evaluation. To construct rubrics for our benchmark, our legal expert annotators marked reference spans (colored text in Figure 1) in the ground-truth answer for each score (Facts, Cited Articles, Analysis). These spans contain the information pertinent to their respective dimension. Each span is then annotated with important span subsets (usually a few words within the span) that are crucial for an answer to be considered correct (bold text in Table 1). Missing this crucial information should result in a lower score. We opted not to assign specific point values or ‘costs’ to each important subset, as previous work with rubrics (Starace et al., 2025; Pires et al., 2025), to minimize the annotation burden. The process involves simply highlighting the three scoring dimensions using different colors and then marking the important subsets within those highlighted sections. The LLM-judge is instructed to determine whether the candidate answer covers the information in the spans and then use this assessment to evaluate each scoring dimension. The complete prompt is presented in Appendix B (Fig. 6).
@@ 5.1 Models
Our experiments evaluate a diverse range of LLMs, comprising proprietary models (from OpenAI, Google, Anthropic) and open-weight models; Deepseek-R1 (DeepSeek-AI et al., 2025), Gemma-3 (Team et al., 2025), and Llama-Krikri-8B (Roussis et al., 2025); a model specifically pretrained for the Greek language. We accessed proprietary models and the large open-weight Deepseek-R1 through Application Programming Interfaces (APIs) provided by OpenAI, Google, and AWS. The remaining open-weight models were deployed on a cluster of eight A100 GPUs using the vLLM framework (Kwon et al., 2023). Due to limited resources, we only evaluate a single run for each model. We used the default parameter configurations as specified by each model’s provider.
@@
To generate responses from LLM candidates, we designed a system and user prompt for the questions of the benchmark. The system prompt instructs the LLM to answer with citations to Greek statutory articles. The user prompt is structured to first describe the overall task, including clear instructions on the expected output format. Then it provides the numbered facts of the case, the question and the relevant legal context. The original prompts are available in Appendix B.
@@ 5.2 Manual Evaluation by Legal Experts
Model Simple-J Span-J Cost

### v3

@@ Abstract
We introduce GreekBarBench, a benchmark that evaluates LLMs on legal questions across five different legal areas from the Greek Bar exams, requiring citations to statutory articles and case facts. To tackle the challenges of free-text evaluation, we propose a three-dimensional scoring system combined with an LLM-as-a-judge approach. We also develop a meta-evaluation benchmark to assess the correlation between LLM-judges and human expert evaluations, revealing that simple, span-based rubrics improve their alignment. Our extensive evaluation of $13$ proprietary and open-weight LLMs shows that even though the top models exhibit impressive performance, they remain susceptible to critical errors, most notably a failure to identify the correct statutory articles.
@@
Odysseas S. Chlapanis1,2Dimitrios Galanis2,3Nikolaos Aletras4Ion Androutsopoulos1,2 1Department of Informatics, Athens University of Economics and Business, Greece 2Archimedes, Athena Research Center, Greece 3Athena Research Center, Greece 4University of Sheffield, United Kingdom
@@ 1 Introduction
As legal AI assistants become increasingly prevalent, the need for realistic legal LLM benchmarks has never been more imperative. [^1] https://www.abajournal.com/web/article/aba-tech-report-finds-that-ai-adoption-is-growing-but-some-are-hesitant Most widely used legal Natural Language Processing (NLP) benchmarks (Chalkidis et al., 2022; Niklaus et al., 2023) focus on classification tasks, e.g., legal judgement prediction (Aletras et al., 2016), which have been criticized (Medvedeva and Mcbride, 2023; Mahari et al., 2023) for being more constrained and less representative than real-world tasks. Even more recent LLM-focused legal benchmarks (Guha et al., 2023; Fei et al., 2024; Joshi et al., 2024) do not go beyond closed-form questions (e.g., multiple-choice questions), failing to capture the true complexity of legal reasoning in practice, which involves identifying, analyzing and synthesizing relevant information to reach a conclusion. Unfortunately, most existing benchmarks with challenging legal questions and free-text responses are proprietary and thus inaccessible to the research community. [^2] https://www.vals.ai/benchmarks
@@
$[2]$ His assistant, Penelope, took a bottle and started washing the area.
$[3]$ The bottle contained pure acetic acid (not a solution, as prescribed), which led to burns on Antonis’ face.
$[4]$ He is now seeking €2.500 for treatment costs plus €75.000 for moral damages.
@@
actions [3][3](Civil Code 922)and Penelope is directly
liable for her negligence [3][3](Civil Code 914).
Hence, both are liable and must compensate Antonis.
Table 1: Cropped example (English translation) from GreekBarBench. The answer requires multi-hop reasoning and citing legal articles and case facts. The spans corresponding to the scoring dimensions are highlighted in color: Facts (green), Cited Articles (blue) and Analysis (orange). Important spans are marked in bold and cited facts are denoted by square brackets. The complete example is presented in Appendix E.
@@
  • A systematic evaluation of 13 frontier and open-weight LLMs on GreekBarBench, using the best LLM-judge at GBB-JME.
  • Manual error analysis to identify important weaknesses in the model responses.
@@ 2 GreekBarBench (GBB)
2.1 Greek Bar Exam
To become licensed attorneys, Greek law graduates must pass the Greek Bar exam, which is organized by the Greek Bar Association and requires a minimum score of six out of ten. The exam evaluates candidates through practical legal questions across five key areas of law: Civil Law, Criminal Law, Commercial Law, Public Law, Lawyers’ Code.
Greece operates under a statutory legal system, where laws are derived from codified statutes rather than judicial precedents. Consequently, the Greek Bar exam is an open-book test requiring candidates to locate and cite the correct statutory articles from the provided legal documents. These documents include the Civil Code and Civil Procedure Code, the Criminal Code and Criminal Procedure Code, eight Commercial Law codes, eleven Public Law codes, the Lawyers’ Code, and the Code of Ethics for Legal Practice (see Table 3).
Past Greek Bar exam papers and suggested solutions (2015–2024) are publicly available in an annually updated PDF booklet. [^4] The booklet is available at https://www.lawspot.gr/nomika-nea/panellinios-diagonismos-ypopsifion-dikigoron-themata-exetaseon-kai-endeiktikes-3 The authors of this paper have been granted permission by the Greek Bar Association to use the questions and by the booklet’s authors to use their respective solutions.
@@ 2.2 Task
Benchmark Lang Cite Legal Judge
Articles Context Eval
@@
Table 2: Comparison of popular legal benchmarks. ‘Lang’: language of dataset. ‘Cite Articles’: the legal articles must be cited. ‘Legal Context’: the necessary legal information is provided. ‘Judge-Eval’: separate benchmark to compare LLMs-as-judges (‘–’ means that the feature is not publicly available).
The design of GreekBarBench is inspired by the workflow of candidate lawyers on the exam. They first study the case facts to identify legal issues, then navigate legal codes to locate the relevant chapter and pinpoint the specific statutory article to support their arguments. To mirror this process, each instance in the benchmark is derived from an exam question and consists of (1) the facts of the incident, (2) the exam question, and (3) a collection of potentially relevant chapters of statutory articles. The desired output is the free-text answer to the question, providing an analysis with citations to the case facts and the applicable statutory articles.
Table 2 highlights the unique contributions of GreekBarBench. It is one of the few benchmarks to require citations to legal articles (and the only one to require citations to case facts). Furthermore, it is unique in providing a corpus of potentially relevant articles and including a dedicated framework for assessing LLMs-as-judges.
@@ 2.3 Dataset Statistics
We collect a total of 65 exam papers; 13 exam papers from each of the five aforementioned areas. The PDF booklet (§ 2.1) is converted to text format and further processed (§ 2.5) to prepare the dataset. Each exam paper includes 4.7 questions on average, resulting in a total of 310 questions. We keep the questions from 2024 (22 in total) as a semi-private test set, to avoid data contamination.555‘Semi-private’ means that the test set is not public, but the raw PDFs might have leaked into the pretraining data. The public test set contains 288 questions, while the semi-private set will be updated annually with two new exam papers from each legal area.
Answering the exam questions requires citing articles from 25 legal code documents, obtained from the official website of the Greek National Printing House. [^6] https://search.et.gr/el/ Detailed statistics for these documents are presented in Table 3 per legal area. Articles are cited 931 times in total, across all exam questions. The articles within each legal code document are grouped thematically into chapters. The total number of citable articles is more than 16,000. This extensive corpus is far too large to fit entirely within the context window of any LLM.
Law Areas Questions Legal Total Cited Context
@@ 2.4 Relevant Legal Context
As mentioned in Section 2.1, the Greek Bar exams are open-book, allowing candidate lawyers to navigate legal code documents to identify relevant statutory articles for the presented case. Simulating this setup presents several challenges. One approach would be implementing a Retrieval-Augmented Generation (RAG) pipeline, using sparse (e.g., BM25) or dense retrievers (Karpukhin et al., 2020) to select the $k$ most ‘relevant’ articles for inclusion in the LLM’s input. However, this approach suffers from three significant limitations: a) candidate lawyers taking the exams do not have access to such retrieval tools, making direct comparisons with human performance problematic; b) retrievers are prone to errors, creating a substantial risk that even with large values of $k$, the ground truth articles might not appear among the top retrieved articles; and c) as demonstrated by Krishna et al. (2025), benchmarking RAG systems requires testing multiple configurations with varying values of $k$ and, ideally, different retriever models, complicating fast integration of new LLMs.
To better simulate the exam environment, we propose an approach that mirrors how a candidate lawyer works: using chapter-level context rather than discrete retrieved articles. For each case, we first use regular expressions to extract the ground truth legal articles cited in its questions. We then identify the legal code chapters these articles belong to and aggregate every article from those chapters into a single corpus. This corpus serves as a broad legal context for all questions related to that case, more faithfully simulating how a lawyer navigates an entire chapter. The average corpus length is 62,000 tokens per question (Table 3), which is within the capacity of most modern LLMs.
@@ 2.5 Fact Segmentation
To streamline the evaluation process, we require citations to facts in candidate answers, though this is not mandatory in the official exams. To implement this, we segment the case facts into sentences using the Segment-Any-Text neural model (Frohmann et al., 2024) and present them as a numbered list (as shown in Table 1). This structure makes it straightforward to detect any factual errors.
@@ 2.6 Three-Dimensional Scoring System
The official evaluation committee of the Greek Bar Exams grades candidate answers on a scale of 1 to 10. This process lacks explicit guidelines and relies on a holistic comparison to the ground truth, limiting its analytical depth for benchmarking LLMs.
Drawing inspiration from established legal research and evaluation practices (Clark and DeSanctis, 2013), and guided by our legal expert annotators (§ 5.2), we develop a novel three-dimensional scoring system to improve the evaluation process for the benchmark. The proposed approach assesses legal reasoning across three dimensions: the Facts, the Cited Articles, and the Analysis. Each dimension is rated on a scale of 1 to 10, and the final score is their mean. This fine-grained framework allows the detection of specific shortcomings in the abilities of LLMs. The Facts score measures understanding of case facts; the Cited Articles score evaluates the ability to identify and cite applicable legal articles; and the Analysis score evaluates the ability to construct valid legal arguments.
@@ 3.1 Simple LLM-Judge
As an initial approach, we designed a straightforward prompt for a simple LLM-judge. The prompt outlines the evaluation task and explicitly defines the criteria for the Facts, Cited Articles, and Analysis scores. All necessary contextual information is provided; the facts of the case, the specific legal question, the ground truth answer with the cited articles and the candidate answer to be evaluated. This context mirrors the information provided to the human annotators for the manual evaluations (§ 5.2). The required output format is clearly specified: the model must provide an explanation for each score, followed by the integer score. The complete prompt is presented in Appendix D (Fig. 4).
@@ 3.2 Span LLM-Judge
Inspired by Clark and DeSanctis (2013), who found that rubrics improve the consistency of legal writing evaluation, we developed a rubric-based annotation process for our benchmark. Our legal experts constructed these rubrics by first identifying reference spans in the ground-truth answer corresponding to the three scoring dimensions, marking each with a distinct color: Facts (green), Cited Articles (blue), and Analysis (orange). An example of these colored spans is shown in Figure 1. Within each span, the annotators then identified important subsets, i.e., the specific words or phrases crucial for a correct answer (shown in bold in Table 1).
To minimize the annotation burden, we did not assign point values to these subsets, which contrasts with the approach in prior work (Starace et al., 2025; Pires et al., 2025). Instead, the LLM-judge is instructed to assess whether a candidate answer covers the key information within the spans, with a focus on the important subsets, and to use this assessment to score each dimension. The full prompt is available in Appendix D (Fig. 5).
@@ 5.1 Models
Our experiments evaluate a diverse range of LLMs, comprising proprietary models (from OpenAI, Google, Anthropic) and open-weight models; Deepseek-R1 (DeepSeek-AI et al., 2025), Gemma-3 (Team et al., 2025), and Llama-Krikri-8B (Roussis et al., 2025); a model specifically pretrained for the Greek language. We accessed proprietary models and the large open-weight Deepseek-R1 through Application Programming Interfaces (APIs) provided by OpenAI, Google, and AWS. The remaining open-weight models were deployed on a cluster of eight A100 GPUs using the vLLM framework (Kwon et al., 2023). Due to limited resources, we performed three runs for each model. We used the default parameter configurations as specified by each model’s provider.
@@
To generate responses from candidate LLMs, we designed a system and user prompt for the benchmark. The system prompt instructs the LLM to answer with citations to Greek statutory articles. The user prompt is structured to first describe the overall task, including clear instructions on the expected output format. Then it provides the numbered facts of the case, the question and the relevant legal context. The original prompt templates are available in Appendix D.
@@ 5.2 Manual Evaluation by Legal Experts
In this section we present the manual evaluations that we collected for GBB-JME (§ 5.3), our Judge Meta-Evaluation benchmark for assessing LLM-judges on GreekBarBench. We obtain ground truth evaluations (Facts, Cited Articles, Analysis scores on a scale of 1 to 10) from two expert legal annotators who are licensed Greek lawyers with practical experience. The annotators were compensated for their time and expertise. They evaluated five LLMs on 87 questions drawn from three exam sessions (2024-A, 2023-A, and 2023-B), resulting in a total of 1,305 annotated samples. The models evaluated on all three exams were Claude-3.7-Sonnet, OpenAI-o1, GPT-4o, and Gemini-2.0-Flash. For the 2024 exam, we included the open-source Llama-3.1-70B; however, due to its poor performance and generation failures on several questions, we replaced it with Deepseek-R1 for the 2023-A and 2023-B exams. Annotations were managed with the open-source platform doccano. [^8] https://doccano.prio.org
As shown in Figure 1, the top models in the 2024 exam were Claude-3.7-Sonnet (7.55) and OpenAI-o1 (7.52). While all LLMs except Llama-3.1-70B passed the exam (score > 6.0), they still performed below the average expert (7.78) and the top 5% of experts (8.87).
5.3 Judge Meta-Evaluation (GBB-JME)
We evaluate seven LLMs-as-judges on our meta-evaluation benchmark, GBB-JME, using two prompts: the Simple-Judge3.1) and the rubric-based Span-Judge3.2). To measure the consistency of the judges, we report the average and standard error over three runs.
The results, presented in Table 4, reveal a clear trend: leading models (GPT-4.1, Gemini-2.0-Flash, and GPT-4.1-mini) generally improve with the Span-Judge prompt while weaker ones seem to struggle with its complexity. Surprisingly, the top-performing model was GPT-4.1-mini (0.862), which outperformed its larger and theoretically more powerful variant, GPT-4.1 (0.837). The stronger performance from the smaller model as a judge corroborates findings from other work (Niklaus et al., 2025).
More generally, the results highlight a key distinction between a model’s performance as a judge and as a candidate, a pattern that holds across different model families. This is illustrated by the open-weight Gemma-3-27B (0.785 with Simple-Judge), which outperformed Gemini-2.0-Flash (0.778 with Span-Judge) despite the latter’s superior performance as a candidate (see §5.4, Table 5).
Due to its strong performance and cost-effectiveness, we adopted GPT-4.1-mini with the Span-Judge prompt for all subsequent evaluations, with the total cost amounting to approximately \$100. For those seeking a more accessible option, the open-weight Gemma-3-27B serves as a strong, low-cost alternative. To support future work, we are publicly releasing the GBB-JME benchmark, enabling other researchers to evaluate new LLMs-as-judges. This approach ensures that the benchmark’s value grows over time, as it can be continuously replicated and validated with increasingly capable models serving as judges.
Model Simple-Judge Span-Judge Cost